adult fickets are 4.50. student tickets are 2.00. the total receipts for 1720 tickets was 5980.00. find the number of adult and student tickets sold.
Answer by mananth(16946)

adult fickets are 4.50. student tickets are 2.00. the total receipts for 1720 tickets was 5980.00. find the number of adult and student tickets sold.
let number of students be x
the number of adults will be 1720-x
2x+4.50*(1720-x)= 5980
2x+7740-4.5x=5980
2x-4.5x=5980-7740
-2.5x= -1760
x= -1760/-2.5
x= 704 students
y= 1720-704= 1016 adults